# A polynomial bound on the pathwidth of graphs edge-coverable by $k$ shortest paths

### v1

@@ 2.2 Path-decompositions
Path-decomposition.
A path-decomposition of a graph $G$ is a sequence $\mathcal{P}=(X_{1},X_{2},\ldots,X_{q})$ of vertex subsets of $G$, called bags, such that for every edge $\{x,y\}\in E$ there is at least one bag containing both endpoints, and for every vertex $x\in V$, the bags containing $x$ form a continuous subsequence of $\mathcal{P}$. The width of $\mathcal{P}$ is $\max\{\lvert X_{i}\rvert-1:1\leq i\leq q\}$, and the pathwidth $\mathrm{pw}(G)$ of $G$ is the smallest width of a decomposition, among all path-decomposition of $G$. The bags $X_{1},X_{q}$ are called the extremities of $\mathcal{P}$. We say that $X_{1}$ is the leftmost bag, and that $X_{q}$ is the rightmost bag.
@@
A layering of a graph $G$ is a partition of its vertices into a sequence $(V_{1},\ldots,V_{t})$ of subsets of $V(G)$ called layers, such that each graph $G[V_{i}]$ is an independent set, and for every $i,j\in\{1,\ldots,t\}$ such that $|j-i|\geq 2$, we have $E_{G}(V_{i},V_{j})=\emptyset$. [^3] Note that our definition of layering differs from the one which is commonly used in the literature, as one usually allows in a layering edges between vertices from a same layer.
A $k$-layering of $G$ is a layering $(V_{1},\ldots,V_{t})$ such that for every $i<t$, we have $\lvert E_{G}(V_{i},V_{i+1})\rvert\leq k$. Note that in particular, it implies that $\lvert N_{V_{i+1}}(V_{i})\rvert\leq k$ and $\lvert N_{V_{i}}(V_{i+1})\rvert\leq k$. The bags $V_{1}$ and $V_{t}$ are the extremities of the layering.
@@
Let $G$ be a graph admitting a $k$-layering. Then $G$ has pathwidth at most $k$. Moreover, if $G$ is connected, then $G$ has a path-decomposition of width at most $k$, whose extremities are exactly the extremities of the $k$-layering.
@@
To prove the “Moreover” part of the lemma, observe that if $G$ is connected, then both $V_{1}$ and $V_{t}$ have size at most $k$. We claim that it is not hard to check that the path-decomposition given by Theorem˜2.2 can then be slightly modified so that its extremities are $V_{1}$ and $V_{k}$. ∎
@@ 3 A General Polynomial Bound
For every $1\leq i\leq k$, we use Theorem˜3.1 to obtain the set $X_{i}$. We consider their union $X:=\bigcup_{i=1}^{k}X_{i}$. Observe that $\lvert X\rvert\leq\sum_{i\in[k]}\lvert X_{i}\rvert\leq k\cdot(720k^{3}+4k)=720k^{4}+4k^{2}$. We now let $C$ be a connected component of GXG-X. As $P_{1},\ldots,P_{k}$ edge-cover (and thus also vertex-cover) $G$, there is some $i\in\{1,\ldots,k\}$ such that the path $P_{i}$ intersects $C$. In particular, as $C$ is included in a connected component of $G-X_{i}$, it implies that $C$ has pathwidth smaller than $6k$, and thus that $\mathrm{pw}(G-X)\leq 6k$. Using Remark˜2.1 we conclude that $\mathrm{pw}(G)\leq\lvert X\rvert+\mathrm{pw}(G-X)\leq 720k^{4}+4k^{2}+6k$. ∎
The remainder of the section will consist in a proof of Theorem˜3.1. In Subsection 3.1 we prove some preliminary results, and show that our proof reduces to the problem of finding some special kind of separations in a subgraph $G_{0}$ of $G$ which is $6k$-layered (see Lemma˜3.9). In Subsection 3.2, we give a proof of Lemma˜3.9, which is the most technical part of our proof.
@@ 3.1 Preliminary Results
We set $U:=\bigcup_{i=1}^{\ell}V(Q_{i})$, $r:=d(a_{1},b_{1})$, and for every $i\in\{0,\ldots,r\}$, we let $U_{i}:=\{u\in U:d(a_{1},u)=i\}$. As every path $Q_{j}$ is parallel to $P_{1}$, observe that for each $(i,j)\in\{0,\ldots,r\}\times\{1,\ldots,\ell\}$, we must have $|V(Q_{j})\cap U_{i}|\leq 1$. In particular, for each $i\in\{0,\ldots,r\}$, we have $|U_{i}|\leq\ell$. Hence $(U_{0},\ldots,U_{r})$ forms a partition of $U$, such that each edge of $G[U]$ has its two endvertices in some pair $(U_{i},U_{i^{\prime}})$ such that $|i^{\prime}-i|\leq 1$. Moreover, as the paths $Q_{j}$ are all parallel to $P_{1}$, no edge of a path $Q_{j}$ can have both endvertices in the same set $U_{i}$, so $(U_{0},\ldots,U_{r})$ is indeed a $\ell$-layering of $\bigcup_{j=1}^{\ell}Q_{j}$. ∎
@@
Let $a\leq_{1}b\leq_{1}c$ be three vertices such that there exists two paths $P,Q$ both parallel to $P_{1}$, such that $P$ connects $a$ to $b$ and $Q$ connects $b$ to $c$. Then $P\cdot Q$ is parallel to $P_{1}$.
@@
Assume that $\mathcal{Q}$ is not reduced, and let $Q$ be a reducing path in $\mathcal{Q}$. We also let $P$ be a subpath of $Q$ with endvertices $u,v$ such that $u<_{1}v$, and $R$ be a shortest path in $G$ from $u$ to $v$ intersecting $P_{1}$. We let $R_{1},R_{2}$ denote two shortest paths respectively from $a_{1}$ to $u$ and from $v$ to $b_{1}$. We also write $Q=Q_{1}\cdot Q_{2}\cdot Q_{3}$, so that $Q_{2}$ is the $uv$-subpath of $Q$ parallel to $P_{1}$ (see Figure˜2). Observe now that both paths $Q^{\prime}:=Q_{1}\cdot R\cdot Q_{3}$ and $Q^{\prime\prime}:=R_{1}\cdot Q_{2}\cdot R_{2}$ are shortest paths in $G$ that both intersect $P_{1}$ and cover all the edges of $Q$. In particular, it implies that the family of paths $\mathcal{Q}^{\prime}:=(\mathcal{Q}\setminus\{Q\})\cup\{Q^{\prime},Q^{\prime\prime}\}$ covers all the edges covered by the paths of $\mathcal{Q}$, and has strictly less paths than $\mathcal{Q}$ that do not intersect $P_{1}$. We apply iteratively the same operation as long as the obtained family $\mathcal{Q}^{\prime}$ is not reduced. In particular, as at each step, the number of paths of $\mathcal{Q}$ which are disjoint with $P_{1}$ strictly decreases, then after at most $|\mathcal{Q}|$ iterations, we obtain a reduced family with the desired properties. ∎
@@
We now partition the set $\mathcal{P}:=\{P_{i}:1\leq i\leq k\}$ into the subset $\mathcal{P}_{1}$ of the paths in $\mathcal{P}$ that intersect $P_{1}$, and $\mathcal{P}_{2}:=\mathcal{P}\setminus\mathcal{P}_{1}$.
@@
We say that the edges of the paths in $\mathcal{A}\cup\mathcal{P}_{2}$ are bad, and we call a vertex $v\in V(G)$ bad if it is the endvertex of a bad edge. We call every vertex which is not bad good, and we let $V_{b},V_{g}$ denote respectively the set of bad and good vertices of $G$. We moreover set $X_{0}:=\{a_{P}:P\in\mathcal{P}_{1}\}\cup\{b_{P}:P\in\mathcal{P}_{1}\}$ and $X_{1}:=\{u_{P}:P\in\mathcal{P}_{1}\}\cup\{v_{P}:P\in\mathcal{P}_{1}\}$. Note that in particular, $|X_{0}|\leq 4k$ and that the only bad vertices that belong to $P_{1}$ are the vertices $a_{P},b_{P}$ for $P\in\mathcal{P}_{1}$, hence $V(P_{1})\cap V_{b}\subseteq X_{0}$.
We now consider the subgraph $G_{0}:=\bigcup_{P\in\mathcal{B}}P$ of $G$. The following remark immediately follows from the definition of $\mathcal{B}$, the fact that $|\mathcal{P}_{1}|\leq|\mathcal{P}|\leq 2k$ and that every path in $\mathcal{B}$ is edge-covered by at most $3$ paths that are parallel to $P_{1}$.
@@
Let $u$ be a good vertex. As $G$ is covered by $\mathcal{P}$, there exists some path $P\in\mathcal{P}$ such that $u\in V(P)$. In particular, as $u$ is a good vertex, we must have $P\in\mathcal{P}_{1}$, and moreover $u$ must be a vertex of $\widetilde{C}_{P}$. It implies that $u\in V(G_{0})$, and thus that $V_{g}\subseteq V(G_{0})$. The ’Moreover’ part follows from the fact that by definition, no bad edge can be incident to a good vertex. In particular, note that every edge from $E(G)\setminus E(G_{0})$ is bad, hence $G[V_{g}]$ is indeed a subgraph of $G_{0}$. ∎
@@
We let $X$ be given by Lemma˜3.8. In particular, by Lemma˜3.7, every component of GXG-X that contains a vertex of $P_{1}$ induces a subgraph of $G_{0}$. By Lemma˜2.3, Lemma˜3.2 and Remark˜3.6, $G_{0}$ is $6k$-layered, and thus has pathwidth at most $6k$, allowing us to conclude the proof of Theorem˜3.1. ∎
@@
We now show that $X$ separates $V(P_{1})$ from $V_{b}$ in $G$. Suppose, for contradiction, that there exists a path $R$ in GXG-X connecting a vertex $v$ from $V_{b}$ to a vertex $u$ on $P_{1}$. Among those paths, choose a path $R$ with minimum length. Since $V(P_{1})\cap V_{b}\subseteq X_{0}\subseteq X$ and $v\notin V(P_{1})$, $R$ has length at least $1$. Also, since $R$ is a path of minimum length connecting $V_{b}$ to $P_{1}$, every vertex in RvR-v must be good. Since every vertex in RvR-v is good, there is no bad edge in $R$. Hence, as $\mathcal{P}$ edge-covers $G$, all the edges of $R$ must be covered by the paths from $\mathcal{B}$, implying that $R$ is also a path of $G_{0}$.
@@
We fix some path $P$ in $G_{0}$ which is parallel to $P_{1}$, and some path $A\in\mathcal{A}\cup\mathcal{P}_{2}$ which intersects $P$. We will in fact prove that one can find a set $X_{A,P}$ of vertices with the desired size that separates in $G_{0}$ all the vertices of $Q$ from $P_{1}$. We let $Q$ denote the minimal subpath of $P$ containing all vertices in $V(A)\cap V(P)$, and let $b,c$ denote the endvertices of $Q$ such that $b\leq_{1}c$. By Remark˜3.6, $G_{0}$ is coverable by at most $6k$ paths which are all parallel to $P_{1}$. In particular, for every $i\geq 0$, the set $V_{i}:=\{v\in V(G_{0}):d_{G}(v,a_{1})=i\}$ contains at most $6k$ vertices, and separates in $G_{0}$ the sets $\bigcup_{j\leq i}V_{j}$ and $\bigcup_{j\geq i}V_{j}$. We also may assume without loss of generality that $Q$ contains at least $2$ vertices, hence $b<_{1}c$ (otherwise, we conclude by choosing $X_{A,P}:=V(Q)=\{b\}$).
For every $v\in V(G_{0})$, we let $\iota(v)$ denote the unique index $i\geq 0$ such that $v\in V_{i}$. For every path $L$ in $G_{0}$ which is parallel to $P_{1}$, if $u,v$ denote the endvertices of $L$ with $u\leq_{1}v$, we define the projection of $L$ on $P_{1}$ as the set $U_{L}:=\{x\in V(P_{1}):\iota(u)<\iota(x)<\iota(v)\}$. We say that $L$ is $\nearrow$-free if there does not exist in $G_{0}$ a path $R$ parallel to $P_{1}$ that connects a vertex $x\in U_{L}$ to a vertex $y\in V(L)$ with $x\leq_{1}y$ (see Figure˜5). Symmetricaly, we say that $L$ is $\searrow$-free if there does not exist in $G_{0}$ a path $R$ parallel to $P_{1}$ that connects a vertex $x\in U_{L}$ to a vertex $y\in V(L)$ with $x\geq_{1}y$. Our main result in this subsection will be Lemma˜3.12, which states that if $Q$ is $\nearrow$-free or $\searrow$-free, then we can separate $Q$ from $P_{1}$ in $G_{0}$ after removing $30k$ vertices. Before proving this, we will first show in the next two lemmas that $Q$ can be covered by at most two subpaths which are each $\nearrow$-free or $\searrow$-free.
@@
If $Q$ is $\nearrow$-free, then we immediately conclude by choosing $Q_{1}$ to be the path having $b$ as a single vertex, thus we may assume that there exists some path $R$ parallel to $P_{1}$ connecting some vertex $x\in U_{Q}$ to a vertex $y\in V(Q)$ such that $x\leq_{1}y$. We moreover choose such a path so that $x$ is maximal with respect to $\leq_{1}$. We let $z$ be the unique vertex of $Q$ such that $\iota(z)=\iota(x)$, and set $Q_{1}:=Q[b,z],Q_{2}:=Q[z,c]$. By maximality of $x$, note that $Q_{2}$ must be $\nearrow$-free. We moreover claim that $Q_{1}$ is $\searrow$-free, as otherwise we would be able to find a shortest path $T$ in $G$ from $b$ to $c$ that intersects $P_{1}$ (see Figure˜7). In particular, as $b,c\in V(A)$ the existence of such a path would imply that $A$ is a reducing path belonging to $\mathcal{P}$, contradicting our assumption that $\mathcal{P}$ is reduced. ∎
Figure 7: The configuration at the end of the proof of Lemma˜3.11. The purple dashed path $R^{\prime}$ represents a shortest path path which is parallel to $P_{1}$, and connects vertices $y^{\prime}\leq_{1}x^{\prime}$, with $y^{\prime}\in V(Q_{1})$ and $x^{\prime}\in U_{Q_{1}}$.
@@
We set $y_{0}:=c_{L}$ and $W_{0}:=V(L)$. We construct a sequence of vertices $y_{\ell}\leq_{1}\cdots\leq_{1}y_{0}$ and a sequence of sets $W_{0},\ldots,W_{\ell}$ for some $0\leq\ell\leq k^{\prime}$. These sequences are constructed by induction as follows. Let $j\geq 0$ and assume that we already constructed $y_{0},\ldots,y_{j}$ and $W_{0},\ldots,W_{j}$ for some $j<k^{\prime}$, such that, up to permuting the indices of the paths $Q_{1},\ldots,Q_{k^{\prime}}$, for every $j^{\prime}\leq j$, $y_{j^{\prime}}\in V(Q_{j^{\prime}})$. We now let $y_{j+1}$ be any vertex which is maximal with respect to $\leq_{1}$ such that
@@
We start showing (i). First, as $y_{j+1}\in W_{j}$, there exists some $j^{\prime}\leq j$ such that $y_{j+1}\in V(Q_{j^{\prime}})$ and $y_{j+1}\leq_{1}y_{j^{\prime}}$. We consider the path $R_{j^{\prime}}$ given by induction hypothesis. In particular, both $y_{j+1}$ and $y_{j^{\prime}}$ belong to $Q_{j^{\prime}}$, hence by Lemma˜3.4, the path $R_{j+1}:=Q_{j^{\prime}}[y_{j+1},y_{j^{\prime}}]\cdot R_{j^{\prime}}$ is parallel to $P_{1}$, and connects $y_{j+1}$ to $c_{L}$ in $G_{0}$. This proves (i).

### v2

@@ 2.2 Path decompositions
Path decomposition.
A path decomposition of a graph $G$ is a sequence $\mathcal{P}=(X_{1},X_{2},\ldots,X_{q})$ of vertex subsets of $G$, called bags, such that $V(G)=\bigcup_{i=1}^{q}X_{i}$, for every edge $\{x,y\}\in E$ there is at least one bag containing both endpoints, and for every vertex $x\in V$, the bags containing $x$ form a continuous subsequence of $\mathcal{P}$. The width of $\mathcal{P}$ is $\max\{\lvert X_{i}\rvert-1:1\leq i\leq q\}$, and the pathwidth $\mathrm{pw}(G)$ of $G$ is the smallest width of a decomposition, among all path decompositions of $G$. The bags $X_{1},X_{q}$ are called the extremities of $\mathcal{P}$. We say that $X_{1}$ is the leftmost bag, and that $X_{q}$ is the rightmost bag.
@@
A layering of a graph $G$ is a partition of its vertices into a sequence $(V_{1},\ldots,V_{t})$ of subsets of $V(G)$ called layers, such that each graph $G[V_{i}]$ is an independent set, and for every $i,j\in\{1,\ldots,t\}$ such that $|j-i|\geq 2$, we have $E_{G}(V_{i},V_{j})=\emptyset$. [^3] Note that our definition of layering differs from the one which is commonly used in the literature, as one usually allows edges between vertices from a same layer.
A $k$-layering of $G$ is a layering $(V_{1},\ldots,V_{t})$ such that for every $0<i<t$, we have $\lvert E_{G}(V_{i},V_{i+1})\rvert\leq k$. Note that in particular, it implies that $\lvert N_{V_{i+1}}(V_{i})\rvert\leq k$ and $\lvert N_{V_{i}}(V_{i+1})\rvert\leq k$. The bags $V_{1}$ and $V_{t}$ are the extremities of the layering.
@@
Let $G$ be a graph admitting a $k$-layering. Then $G$ has pathwidth at most $k$. Moreover, if $G$ is connected, then $G$ has a path decomposition of width at most $k$, whose extremities are exactly the extremities of the $k$-layering.
@@
To prove the “Moreover” part of the lemma, observe that if $G$ is connected, then both $V_{1}$ and $V_{t}$ have size at most $k$. We claim that it is not hard to check that the path decomposition given by Theorem˜2.2 can then be slightly modified so that its extremities are $V_{1}$ and $V_{k}$. ∎
@@ 3 A General Polynomial Bound
For every $1\leq i\leq k$, we use Theorem˜3.1 to obtain the set $X_{i}$. We consider their union $X:=\bigcup_{i=1}^{k}X_{i}$. Observe that $\lvert X\rvert\leq\sum_{i\in[k]}\lvert X_{i}\rvert\leq k\cdot(720k^{3}+4k)=720k^{4}+4k^{2}$. We now let $C$ be a connected component of GXG-X. As $P_{1},\ldots,P_{k}$ edge-cover (and thus also vertex-cover) $G$, there is some $i\in\{1,\ldots,k\}$ such that the path $P_{i}$ intersects $C$. In particular, as $C$ is included in a connected component of $G-X_{i}$, it implies that $C$ has pathwidth at most $6k$, and thus that $\mathrm{pw}(G-X)\leq 6k$. Using Remark˜2.1 we conclude that $\mathrm{pw}(G)\leq\lvert X\rvert+\mathrm{pw}(G-X)\leq 720k^{4}+4k^{2}+6k$. ∎
The remainder of the section will consist in a proof of Theorem˜3.1. In Section 3.1 we prove some preliminary results, and show that our proof reduces to the problem of finding a special kind of separations in a subgraph $G_{0}$ of $G$ which is $6k$-layered (see Lemma˜3.9). In Section 3.2, we give a proof of Lemma˜3.9, which is the most technical part of our proof.
@@ 3.1 Preliminary Results
We set $U:=\bigcup_{i=1}^{\ell}V(Q_{i})$, $r:=d(a_{1},b_{1})$, and for every $i\in\{0,\ldots,r\}$, we let $U_{i}:=\{u\in U:d(a_{1},u)=i\}$. As every path $Q_{j}$ is parallel to $P_{1}$, observe that for each $(i,j)\in\{0,\ldots,r\}\times\{1,\ldots,\ell\}$, we must have $|V(Q_{j})\cap U_{i}|\leq 1$. In particular, for each $i\in\{0,\ldots,r\}$, we have $|U_{i}|\leq\ell$. Hence, $(U_{0},\ldots,U_{r})$ forms a partition of $U$, such that each edge of $G[U]$ has its two endvertices in some pair $(U_{i},U_{i^{\prime}})$ such that $|i^{\prime}-i|\leq 1$. Moreover, as the paths $Q_{j}$ are all parallel to $P_{1}$, no edge of a path $Q_{j}$ can have both endvertices in the same set $U_{i}$, so $(U_{0},\ldots,U_{r})$ is indeed a $\ell$-layering of $\bigcup_{j=1}^{\ell}Q_{j}$. ∎
@@
Let $a\leq_{1}b\leq_{1}c$ be three vertices such that there exist two paths $P,Q$ both parallel to $P_{1}$, such that $P$ connects $a$ to $b$ and $Q$ connects $b$ to $c$. Then $P\cdot Q$ is parallel to $P_{1}$.
@@
Assume that $\mathcal{Q}$ is not reduced, and let $Q$ be a reducing path in $\mathcal{Q}$. We also let $P$ be a subpath of $Q$ with endvertices $u,v$ such that $u<_{1}v$, and $R$ be a shortest path in $G$ from $u$ to $v$ intersecting $P_{1}$. We let $R_{1},R_{2}$ denote two shortest paths respectively from $a_{1}$ to $u$ and from $v$ to $b_{1}$. We also write $Q=Q_{1}\cdot Q_{2}\cdot Q_{3}$, so that $Q_{2}$ is the $uv$-subpath of $Q$ parallel to $P_{1}$ (see Figure˜2). Observe now that both paths $Q^{\prime}:=Q_{1}\cdot R\cdot Q_{3}$ and $Q^{\prime\prime}:=R_{1}\cdot Q_{2}\cdot R_{2}$ are shortest paths in $G$ that both intersect $P_{1}$ and cover all the edges of $Q$. In particular, it implies that the family of paths $\mathcal{Q}^{\prime}:=(\mathcal{Q}\setminus\{Q\})\cup\{Q^{\prime},Q^{\prime\prime}\}$ covers all the edges covered by the paths of $\mathcal{Q}$, and has strictly fewer paths than $\mathcal{Q}$ that do not intersect $P_{1}$. We apply iteratively the same operation as long as the obtained family $\mathcal{Q}^{\prime}$ is not reduced. In particular, as at each step, the number of paths of $\mathcal{Q}$ which are disjoint with $P_{1}$ strictly decreases, then after at most $|\mathcal{Q}|$ iterations, we obtain a reduced family with the desired properties. ∎
@@
We now partition the set $\mathcal{P}:=\{P_{i}:1\leq i\leq 2k\}$ into the subset $\mathcal{P}_{1}$ of the paths in $\mathcal{P}$ that intersect $P_{1}$, and $\mathcal{P}_{2}:=\mathcal{P}\setminus\mathcal{P}_{1}$.
@@
We say that the edges of the paths in $\mathcal{A}\cup\mathcal{P}_{2}$ are bad, and we call a vertex $v\in V(G)$ bad if it is the endvertex of a bad edge. We call every vertex which is not bad good, and we let $V_{b}$ and $V_{g}$ denote respectively the set of bad vertices and the set of good vertices of $G$. We moreover set $X_{0}:=\{a_{P}:P\in\mathcal{P}_{1}\}\cup\{b_{P}:P\in\mathcal{P}_{1}\}$ and $X_{1}:=\{u_{P}:P\in\mathcal{P}_{1}\}\cup\{v_{P}:P\in\mathcal{P}_{1}\}$. Note that in particular, $|X_{0}|\leq 4k$ and that the only bad vertices that might belong to $P_{1}$ are the vertices $a_{P},b_{P}$ for $P\in\mathcal{P}_{1}$ whenever they are respectively equal to $u_{P},v_{P}$, hence $V(P_{1})\cap V_{b}\subseteq X_{0}$.
We now consider the subgraph $G_{0}:=\bigcup_{P\in\mathcal{B}}P$ of $G$. The following remark immediately follows from the definition of $\mathcal{B}$, the fact that $|\mathcal{P}_{1}|\leq|\mathcal{P}|\leq 2k$, and that every path in $\mathcal{B}$ is edge-covered by at most $3$ paths that are parallel to $P_{1}$.
@@
Let $u$ be a good vertex. As $G$ is covered by $\mathcal{P}$, there exists some path $P\in\mathcal{P}$ such that $u\in V(P)$. In particular, as $u$ is a good vertex, we must have $P\in\mathcal{P}_{1}$, and moreover $u$ must be a vertex of $\widetilde{C}_{P}$. It implies that $u\in V(G_{0})$, and thus that $V_{g}\subseteq V(G_{0})$. The second part of the lemma follows from the fact that by definition, no bad edge can be incident to a good vertex. In particular, note that every edge from $E(G)\setminus E(G_{0})$ is bad, hence $G[V_{g}]$ is indeed a subgraph of $G_{0}$. ∎
@@
We let $X$ be given by Lemma˜3.8. In particular, by Lemma˜3.7, every component of GXG-X that contains a vertex of $P_{1}$ induces a subgraph of $G_{0}$. By Lemma˜2.3, Lemma˜3.2 and Remark˜3.6, $G_{0}$ is $6k$-layered, and thus has pathwidth at most $6k$, allowing us to conclude the proof of Theorem˜3.1. By Remark˜3.6, $G_{0}$ is edge-covered by at most $6k$ paths parallel to $P_{1}$, which by Lemma˜3.4 implies that it has a $6k$-layering, and thus by Lemma˜2.3, its pathwidth is at most $6k$. ∎
@@
We now show that $X$ separates $V(P_{1})$ from $V_{b}$ in $G$. Suppose, for contradiction, that there exists a path $R$ in GXG-X connecting a vertex $v$ from $V_{b}$ to a vertex $u$ on $P_{1}$. Among those paths, choose a path $R$ with minimum length. Since $V(P_{1})\cap V_{b}\subseteq X_{0}\subseteq X$, we have $v\notin V(P_{1})$, thus $R$ has length at least $1$. Also, since $R$ is a path of minimum length connecting $V_{b}$ to $P_{1}$, every vertex in RvR-v must be good. Since every vertex in RvR-v is good, there is no bad edge in $R$. Hence, as $\mathcal{P}$ edge-covers $G$, all the edges of $R$ must be covered by the paths from $\mathcal{B}$, implying that $R$ is also a path of $G_{0}$.
@@
We fix some path $P$ in $G_{0}$ which is parallel to $P_{1}$, and some path $A\in\mathcal{A}\cup\mathcal{P}_{2}$ which intersects $P$. We let $Q$ denote the minimal subpath of $P$ containing all vertices in $V(A)\cap V(P)$, and let $b,c$ denote the endvertices of $Q$ such that $b\leq_{1}c$. We will in fact prove that one can find a set $X_{A,P}$ of vertices with the desired size that separates in $G_{0}$ all the vertices of $Q$ from $P_{1}$. By Remark˜3.6, $G_{0}$ is coverable by at most $6k$ paths which are all parallel to $P_{1}$. In particular, for every $i\geq 0$, the set $V_{i}:=\{v\in V(G_{0}):d_{G}(v,a_{1})=i\}$ contains at most $6k$ vertices, and separates in $G_{0}$ the sets $\bigcup_{j\leq i}V_{j}$ and $\bigcup_{j\geq i}V_{j}$. We also may assume without loss of generality that $Q$ contains at least $2$ vertices, hence $b<_{1}c$ (otherwise, we conclude by choosing $X_{A,P}:=V(Q)=\{b\}$).
For every $v\in V(G_{0})$, we let $\iota(v)$ denote the unique index $i\geq 0$ such that $v\in V_{i}$. For every path $L$ in $G_{0}$ which is parallel to $P_{1}$, if $u,v$ denote the endvertices of $L$ with $u\leq_{1}v$, we define the projection of $L$ on $P_{1}$ as the set $U_{L}:=\{x\in V(P_{1}):\iota(u)<\iota(x)<\iota(v)\}$. We say that $L$ is $\nearrow$-free if there does not exist in $G_{0}$ a path $R$ parallel to $P_{1}$ that connects a vertex $x\in U_{L}$ to a vertex $y\in V(L)$ with $x\leq_{1}y$ (see Figure˜5). Symmetrically, we say that $L$ is $\searrow$-free if there does not exist in $G_{0}$ a path $R$ parallel to $P_{1}$ that connects a vertex $x\in U_{L}$ to a vertex $y\in V(L)$ with $x\geq_{1}y$. Our main result in this subsection will be Lemma˜3.12, which states that if $Q$ is $\nearrow$-free or $\searrow$-free, then we can separate $Q$ from $P_{1}$ in $G_{0}$ after removing $30k$ vertices. Before proving this, we will first show in the next two lemmas that $Q$ can be covered by at most two subpaths which are each $\nearrow$-free or $\searrow$-free.
@@
If $Q$ is $\nearrow$-free, then we immediately conclude by choosing $Q_{1}$ to be the path having $b$ as a single vertex, thus we may assume that there exists some path $R$ parallel to $P_{1}$ connecting some vertex $x\in U_{Q}$ to a vertex $y\in V(Q)$ such that $x\leq_{1}y$. We moreover choose such a path so that $x$ is maximal with respect to $\leq_{1}$. We let $z$ be the unique vertex of $Q$ such that $\iota(z)=\iota(x)$, and set $Q_{1}:=Q[b,z],Q_{2}:=Q[z,c]$. By maximality of $x$, note that $Q_{2}$ must be $\nearrow$-free. We moreover claim that $Q_{1}$ is $\searrow$-free. Assume for a contradiction that it is not the case, and that there exists a shortest path $R^{\prime}$ parallel to $P_{1}$ connecting a vertices $y^{\prime}\leq_{1}x^{\prime}$, with $y^{\prime}\in V(Q_{1})$ and $x^{\prime}\in U_{Q_{1}}$ (see Figure˜7). Then the path $T:=Q_{1}[b,y^{\prime}]\cdot R^{\prime}\cdot P_{1}[x^{\prime},x]\cdot R\cdot Q_{2}[y,c]$ is a path in $G$ from $b$ to $c$ that intersects $P_{1}$ and which is parallel to $P_{1}$. In particular, as $b,c\in V(A)$ the existence of $T$ implies that $A$ is a reducing path belonging to $\mathcal{P}$, contradicting our assumption that $\mathcal{P}$ is reduced. ∎
Figure 7: The configuration at the end of the proof of Lemma˜3.11. The purple dashed path $R^{\prime}$ represents a shortest path which is parallel to $P_{1}$, and connects vertices $y^{\prime}\leq_{1}x^{\prime}$, with $y^{\prime}\in V(Q_{1})$ and $x^{\prime}\in U_{Q_{1}}$.
@@
Formally, we set $y_{0}:=c_{L}$ and $W_{0}:=V(L)$. We construct a sequence of vertices $y_{\ell}\leq_{1}\cdots\leq_{1}y_{0}$ and a sequence of sets $W_{0},\ldots,W_{\ell}$ for some $0\leq\ell\leq k^{\prime}$. These sequences are constructed by induction as follows. Let $j\geq 0$ and assume that we already constructed $y_{0},\ldots,y_{j}$ and $W_{0},\ldots,W_{j}$ for some $j<k^{\prime}$, such that, up to permuting the indices of the paths $Q_{1},\ldots,Q_{k^{\prime}}$, for every $j^{\prime}\leq j$, $y_{j^{\prime}}\in V(Q_{j^{\prime}})$. We now let $y_{j+1}$ be any vertex which is maximal with respect to $\leq_{1}$ such that
@@
We start showing (i). First, as $y_{j+1}\in W_{j}$, there exists some $0\leq j^{\prime}\leq j$ such that $y_{j+1}\in V(Q_{j^{\prime}})$ and $y_{j+1}\leq_{1}y_{j^{\prime}}$. We consider the path $R_{j^{\prime}}$ given by induction hypothesis. In particular, both $y_{j+1}$ and $y_{j^{\prime}}$ belong to $Q_{j^{\prime}}$, hence by Lemma˜3.4, the path $R_{j+1}:=Q_{j^{\prime}}[y_{j+1},y_{j^{\prime}}]\cdot R_{j^{\prime}}$ is parallel to $P_{1}$, and connects $y_{j+1}$ to $c_{L}$ in $G_{0}$. This proves (i).